OpenStudy (anonymous):

f(x)=x^2+2x-1 Prove that it's differentiable in [0,1] Using: Lim h-->0+ (f(a+h)-f(a))/h = Lim h--> 0- (f(a+h)-f(a))/h

OpenStudy (anonymous):

Prove that there exists a derivative and that its equivalent on both sides, throughout that interval.

OpenStudy (anonymous):

*it's

OpenStudy (anonymous):

How would you prove the left and right hand derivatives are equivalent on both sides throughout the interval specifically?

OpenStudy (anonymous):

Note that: \[ \lim_{h\to 0^+}\frac{f(x+h)-f(x)}{h}= \lim_{h\to 0^+}\frac{(x+h)^2+2(x+h)-1-(2x^2+2x-1)}{h} \]Show that this is equal to: \[ \lim_{h\to 0^+}\frac{(x+h)^2+2(x+h)-1-(2x^2+2x-1)}{h}=2x+2 \]And, that: \[ \lim_{h\to 0^-}\frac{(x+h)^2+2(x+h)-1-(2x^2+2x-1)}{h}=2x+2 \]And then you have proved it for all real numbers. Clearly \([0,1]\subset\mathbb{R}\). So this holds.

OpenStudy (anonymous):

Ok I get it! Because h--> 0 from both sides is just 0 since it's so close it's negligible?

OpenStudy (anonymous):

By the way, is y=x^(2/3) a polynomial?

OpenStudy (anonymous):

Essentially, but it's not always true of all functions. And, no, it is not.